Find the slope-intercept form for line line that satisfy the conditions given. slope 20, passing through (3,-5).
Answer by srinivas.g(540)

slope-intercept form y= mx+c
where m =slope
c= constant
slope (m)=20
so y= 20x+c..........eq(1)
but this line is passing through (3,-5)
hence
-5 =20*3+c
-5 = 60 +c
move 60 to the left
-5-60 =c
-65 =c
put c= -65 in eq(1)
y=20x-65
result : slope intercept form of line we are looking for is y=20 x -65